OpenStudy (anonymous):

Find the domain, range and two vertical asymptotes of the following function:f(x) = - 2 cot 3x

jimthompson5910 (jim_thompson5910):

cot(3x) = cos(3x)/sin(3x)

jimthompson5910 (jim_thompson5910):

the denominator cannot be zero, if it were then sin(3x) = 0 3x = arcsin(0) 3x = pi*n ... n is any integer x = (pi/3)*n

jimthompson5910 (jim_thompson5910):

this means that domain is basically the set of all real numbers...but...you must kick out values of the form x = (pi/3)*n, n is any integer, to avoid division by zero

jimthompson5910 (jim_thompson5910):

also, this means that the vertical asymptotes are x = (pi/3)*n or x = pi/3, x = 2pi/3, x = pi, etc etc

jimthompson5910 (jim_thompson5910):

with tangent and cotangent, the range is always the set of real numbers so that applies to -2*cot(3x) as well

OpenStudy (anonymous):

Thank you! c:

jimthompson5910 (jim_thompson5910):

you're welcome